Let y1 and y2 be two solutions of the equation a2(x)y'' + a1(x)y' + a0(x)y =0. If W(y1, y2) is the Wronskian of y1 and y2, show that. a2(x)(dw/dx) + a1(x)W = 0
1
Expert's answer
2020-03-06T15:45:41-0500

"W=\begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}=y_1y_2'-y_2y_1'" and "W'=(y_1y_2'-y_2y_1')'=y_1y_2''-y_2y_1''"

"a_2W'+a_1W=a_2(y_1y_2''-y_2y_1'')+a_1(y_1y_2'-y_2y_1')="

"=y_1(a_2y_2''+a_1y_2')-y_2(a_2y_1''+a_1y_1')=y_1(-a_0y_2)-y_2(-a_0y_1)=0"